# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v4

@@ VI-B Embodied Task Planning
Traditional embodied task planning methods are often based on explicit rules and logical reasoning. For example, symbolic planning algorithms such as STRIPS [313] and PDDL[314], and search algorithms like MCTS [315] and A* [316], are used to generate plans. However, these methods often rely on predefined rules, constraints, and heuristics that are rigid and may not adapt well to dynamic or unforeseen changes in the environment. With the popularity of LLMs trained on large-scale data, many researchers have recently attempted to use LLMs for planning or to combine traditional methods with LLMs, leveraging the rich world knowledge embedded within them for reasoning and planning without the need for handcrafted definitions, greatly enhancing the model’s generalization capabilities.
@@ VI-B1 Planning utilizing the Emergent Capabilities of LLMs
Before the scale-up of natural language models, task planners were similarly implemented by training models like BERT on embodied instruction datasets such as Alfred [317] and Alfworld [318], as demonstrated by FILM [319]. However, this approach was limited by the examples in the training set and could not effectively align with the physical world. Nowadays, thanks to the emergent capabilities of large language models (LLMs), LLMs can decompose abstract tasks using their internal world knowledge and chain-of-thought reasoning, similar to how humans reason through task completion steps before acting. For example, Translated LM [320] and Inner Monologue [321] can break down complex tasks into manageable steps and devise solutions using their internal logic and knowledge systems without additional training.
Additionally, some approaches abstract past successful examples into a series of skills stored in a memory bank to consider during inference and improve planning success rates [322, 323, 324] . Interestingly, improving chain-of-thought (CoT) reasoning [325] is also a viable approach. ReAct [326] incorporates CoT reasoning with plan generation, providing more logical action generation. And some works utilize code as the reasoning medium instead of natural language. The Chain of Code method is particularly suitable for embodied agents in reasoning and executing tasks, where task planning is generated as code based on the available API library[327, 328, 329]. Furthermore, multi-turn reasoning can effectively correct potential hallucinations in task planning, a focus of many LLM-based agent studies. For instance, Socratic Models [330] and Socratic Planner [331] use Socratic questioning to derive reliable planning.
However, if LLM is considered as the plan generator since LLM’s plan generation is based on token probability distribution rather than logical inference, it cannot ensure that the plans it generates are logically correct. Therefore, some works consider LLMs as world models rather than plan generators, using MCTS[315] for plan sequence search [332, 333, 334]. Another line of work considers LLMs as instruction translators, leveraging their powerful semantic understanding capabilities to translate plans into planning-specific languages such as PDDL, than leveraging traditional AI planners for planning [335, 336, 337, 338].
During task planning, potential failures that may occur during execution must be considered. These failures often result from the planner not fully accounting for the complexity of the real environment and the difficulty of task execution [321, 339]. Multi-agent collaboration framework ReAd [340] is proposed for efficient self-refinement of plans. Due to a lack of visual information, planned subtasks may deviate from the actual scenario, leading to task failure. Therefore, integrating visual information into planning or replanning during execution is necessary. This approach can significantly enhance the accuracy and feasibility of task planning, better addressing the challenges of real-world environments.
@@ VI-B2 Planning utilizing the visual information from embodied perception model
Based on the above discussion, it is particularly important to further integrate visual information into task planning (or replanning). In this process, object labels, locations, or descriptions provided by visual input can offer critical references for task decomposition and execution by large language models (LLMs). For instance, through visual information, LLMs can more accurately identify target objects and obstacles in the current environment, thereby optimizing task steps or modifying subtask objectives. Some works use an object detector to query the objects present in the environment during task execution and feed this information back to the LLM, allowing it to modify unreasonable steps in the current plan [339, 330, 341]. RoboGPT considers the different names of similar objects within the same task, further improving the feasibility of replanning [12]. However, the information provided by labels is still too limited. Can further scene information be provided? SayPlan [342] proposes using hierarchical 3D scene graphs to represent the environment, effectively mitigating the challenges of task planning in large, multi-floor, and multi-room settings. Similarly, ConceptGraphs [343] also adopts 3D scene graphs to provide environmental information to LLMs. Compared to SayPlan, it offers more detailed open-world object detection and presents task planning in a code-based format, which is more efficient and better suited to the demands of complex tasks.
However, limited visual information may lead to insufficient environmental understanding by the agent. Although visual input can provide important environmental cues, its limitation lies in the inability to capture the complexity and dynamic changes of the environment fully. Without comprehensive visual information, the agent may misinterpret the environmental conditions, resulting in task failure. For instance, if a towel is locked in the bathroom cabinet, the agent may repeatedly search the bathroom without realizing this possibility [12]. Therefore, it is necessary to develop more robust algorithms that integrate multiple sensory data to compensate for the shortcomings of visual information. By incorporating multimodal perception technologies, such as vision, touch, and hearing, the agent’s understanding of the environment can be enhanced. Additionally, leveraging historical data and contextual reasoning can assist the agent in making reasonable judgments and decisions even with limited visual information. This planning approach, combining multimodal fusion and contextual reasoning, not only increases the success rate of task execution but also offers new perspectives for the further development of embodied AI.
@@ VI-B3 Planning utilizing the VLMs
Compared to converting environmental information into text using external visual models, VLM models can capture visual details in latent space, particularly contextual information that is difficult to represent with object labels. VLMs can discern rules underlying visual phenomena; for instance, even if a towel is not visible in the environment, it can be inferred that the towel might be stored in a cabinet. This process essentially demonstrates how abstract visual features and structured textual features can be more effectively aligned in latent space.
In EmbodiedGPT [307], the Embodied-Former module aligns embodied, visual, and textual information, effectively considering the agent’s state and environmental information during task planning. Unlike EmbodiedGPT, which directly uses third-person perspective images, LEO [344] encodes 2D egocentric images and 3D scenes into visual tokens. This method effectively perceives 3D world information and executes tasks accordingly. Similarly, the EIF-Unknow model utilizes Semantic Feature Maps extracted from Voxel Features as visual tokens, which are input along with text tokens into a trained LLaVA model for task planning [345]. Furthermore, embodied multimodal foundation models, or VLA models, have been extensively trained with large datasets in studies like the RT series [13, 305] and PaLM-E [304] to achieve alignment of visual and textual features in embodied scenarios. Additionally, innovative research such as Matcha [346] explores the integration of LLMs with multimodal perception modules (e.g., weight, touch, sound) to enhance robot planning capabilities. For example, when executing the task “pick up a plastic block,” the robot can assess its weight by lifting it, determine if the sound it makes when tapped matches expected plastic sounds, or use tactile feedback to sense the block’s hardness or softness. Moreover, the VLP [347] model enhances planning by considering historical information and predicting future visual states. Once the LLM generates an action, the video model can simulate multiple potential video representations for each action. These video models then serve as heuristic functions to evaluate the proximity of each action to the goal, thereby improving planning effectiveness.
@@ VI-C1 Action utilizing APIs
This hierarchical planning paradigm allows agents to focus more on high-level task planning and decision-making while delegating specific action execution to policy models, simplifying the agent’s development process. Additionally, this modularization of task planners and action planners enables independent development, testing, and optimization, thereby enhancing system flexibility and maintainability. This approach allows agents to adapt to various tasks and environments by invoking different action planners and facilitating modifications to the action planner without significant changes to the agent’s structure. However, invoking external policy models may introduce latency, particularly in real-time tasks, potentially affecting system response speed and efficiency. The most critical issue to address is that the agent’s performance is highly dependent on the quality of the policy model. If the policy model fails to solve the tasks effectively, the overall performance of the agent will be compromised.
@@ VII-A1 Generation-based Methods
LEGENT[363]) 2024 3D Environments Generation
@@ VII-A2 Prediction-based Methods
The prediction-based world model predicts and understands the environment by constructing and utilizing internal representations. By reconstructing corresponding features in the latent space based on provided conditions, it captures deeper semantics and associated world knowledge. This model maps input information to a latent space and operates within that space to extract and utilize high-level semantic information, thereby enabling the robots to perceive the essential representation of the world environment (e.g., I-JEPA[18], MC-JEPA[354], A-JEPA[355], Point-JEPA [364], IWM[356]) and more accurately perform embodied downstream tasks (e.g., iVideoGPT[357]), STP[358], MuDreamer[359]). Compared to pixel-level information, features in the latent space can abstract and decouple various forms of knowledge, allowing the model to handle complex tasks and scenes more effectively and improve its generalization capability[365]. For instance, in spatiotemporal modeling, the world model needs to predict the post-interaction state of an object based on its current state and the nature of the interaction, combining this information with its internal knowledge. Specifically, an embodied world model generates dynamic predictions of the environment by integrating perceptual information and prior knowledge. This approach relies not only on sensory data but also on inherent world knowledge to infer and predict environmental changes, thereby producing more accurate spatiotemporal predictions[357, 358, 359]. This process considers both the current state of objects and their historical data and contextual information.
@@ VII-A3 Knowledge-driven Methods
Knowledge-driven world models inject artificially constructed knowledge into the models, endowing them with world knowledge. This method has shown broad application potential in the field of embodied AI. For example, in the real2sim2real approach[367], real-world knowledge is used to build physics-compliant simulators, which are then used to train robots, enhancing model robustness and generalization capabilities. Additionally, artificially constructing common sense or physics-compliant knowledge and applying them to generative models or simulators is a common strategy (e.g., ElastoGen [361], One-2-3-45 [362], PLoT [360]). This approach imposes more physically accurate constraints on the model, enhancing its reliability and interpretability in generation tasks. These constraints ensure the model’s knowledge is both accurate and consistent, reducing uncertainty during training and application. Some approaches combine artificially created physical rules with LLMs or MLMs. By leveraging the commonsense capabilities of LLMs and LMMs, these approaches (e.g., Holodeck[73], LEGENT[363]) generate diverse and semantically rich scenes through automatic spatial layout optimization. This greatly advances the development of general-purpose embodied agents by training them in novel and diverse environments.
@@ VII-B Data Collection and Training
Figure 16: The illustration of demonstration data collection. The yellow box on the left features operational demonstrations of the Franka and WidowX robotic arms, while human demonstrations are shown in the blue box. On the right, the yellow box showcases operational scenarios of the UR5e and Franka robotic arms in simulation environments, and the blue box displays labeled simulation data. The yellow box at the bottom presents the data formats of these datasets.
@@ VII-C Embodied Control
Figure 18: Examples of embodied control for various locomotion modes, demonstrating the robots’ agile movement and interaction capabilities. The orange box on the left showcases the locomotion of two quadruped robots (Unitree B1, ANYmal) and one biped robot (ATRIAS). The red box on the right displays motion scenes of the Atlas Robot climbing stairs, Unitree H1 waving and high-fiving, and the Boston robot leaping.
@@
Embodied control integrates RL and sim-to-real techniques to optimize strategies through environmental interaction, enabling the exploration of unknown domains, potentially surpassing human capabilities, and adapting to unstructured environments. While robots can mimic many human behaviors, effective task completion often requires RL training based on environmental feedback. The most challenging scenarios include contact-intensive tasks, where manipulation necessitates real-time adjustments based on feedback such as the state, deformation, material, and force of manipulated objects. In such cases, RL is indispensable. In the era of MLMs, these models possess a generalized understanding of scene semantics, providing robust reward functions for RL. Additionally, RL is crucial for aligning large models with their intended tasks [397]. In the future, after pre-training and fine-tuning, RL is still required to align with the physical world, ensuring effective deployment in real-world environments.
@@ VIII Challenges and Future Directions
Efficient Utilization of Human Demonstration Data: Efficient utilization of human demonstration data involves leveraging the actions and behaviors demonstrated by humans to train and improve robotic systems. This process includes collecting, processing, and learning from large-scale, high-quality datasets where humans perform tasks that robots are intended to learn. Current work R3M [398] used action labels and human demonstration data to learn generalizable representations has shown high success rates in some robot grasping tasks, but the efficiency for complex tasks still needs improvement. Therefore, it is important to effectively utilize large amounts of unstructured, multi-label, and multi-modal human demonstration data combined with action label data to train embodied models that can learn various tasks in relatively short periods. By efficiently utilizing human demonstration data, robotic systems can achieve higher levels of performance and adaptability, making them more capable of performing complex tasks in dynamic environments.
@@
Long-Horizon Task Execution: Executing single instructions can often entail long-horizon tasks for robots, exemplified by commands like “clean the kitchen,” which involve activities such as rearranging objects, sweeping floors, wiping tables, and more. Accomplishing such tasks successfully necessitates the robot’s ability to plan and execute a sequence of low-level actions over extended time spans. While current high-level task planners have shown initial success, they often prove inadequate in diverse scenarios due to their lack of tuning for embodied tasks. Addressing this challenge requires the development of efficient planners equipped with robust perception capabilities and much commonsense knowledge.
Causal Relation Discovery: Existing data-driven embodied agents make decisions based on the intrinsic correlations within the data. However, this modeling approach does not allow the models to truly understand the causal relations between knowledge, behavior, and environment, resulting in biased strategies. This makes it difficult to ensure that they can operate in real-world environments in an interpretable, robust, and reliable manner. Therefore, it is important for embodied agents to be driven by world knowledge, capable of autonomous causal reasoning. By understanding the world through interaction and learning its workings via abductive reasoning [399], we can further enhance the adaptability, decision reliability, and generalization capabilities of multimodal embodied agents in complex real-world environments.
For embodied tasks, it is necessary to introduce causal representation learning. This involves establishing spatial-temporal causal relations across modalities through interactive instructions and state predictions, forming a representation learning system based on interaction and deduction. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [400], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied perception, reasoning, and interaction.

### v5

@@ VI-B Embodied Task Planning
Traditional embodied task planning methods are usually based on explicit rules and logical reasoning. For example, symbolic planning algorithms such as STRIPS [309] and PDDL[310], and search algorithms like MCTS [311] and A* [312], are used to generate plans. However, these methods often rely on predefined rules, constraints, and heuristics that are rigid and may not adapt well to dynamic or unforeseen changes in the environment. With the popularity of LLMs, many works have attempted to use LLMs for planning or to combine traditional methods with LLMs, leveraging the rich world knowledge embedded within them for reasoning and planning without the need for handcrafted definitions, greatly enhancing the model’s generalization capabilities.
@@ VI-B1 Planning utilizing the Emergent Capabilities of LLMs
Before the scale-up of natural language models, task planners were similarly implemented by training models like BERT on embodied instruction datasets such as Alfred [313] and Alfworld [314], as demonstrated by FILM [315]. However, this approach was limited by the examples in the training set and could not effectively align with the physical world. Nowadays, thanks to the emergent capabilities of LLMs, they can decompose abstract tasks using their internal world knowledge and chain-of-thought reasoning, similar to how humans reason through task completion steps before acting. For example, Translated LM [316] and Inner Monologue [317] can break down complex tasks into manageable steps and devise solutions using their internal logic and knowledge systems without additional training, such as ReAct [318]. Similarly, the multi-agent collaboration framework ReAd [319] is proposed for efficient self-refinement of plans via the different prompts. Additionally, some approaches abstract past successful examples into a series of skills stored in a memory bank to consider during inference and improve planning success rates [320, 321, 322]. And some works utilize code as the reasoning medium instead of natural language. where task planning is generated as code based on the available API library[323, 324, 325]. Furthermore, multi-turn reasoning can effectively correct potential hallucinations in task planning, a focus of many LLM-based agent studies. For instance, Socratic Models [326] and Socratic Planner [327] use Socratic questioning to derive reliable planning.
However, during task planning, potential failures may occur during execution, often resulting from the planner not fully accounting for the complexity of the real environment and the difficulty of task execution [317, 328]. Due to a lack of visual information, planned subtasks may deviate from the actual scenario, leading to task failure. Therefore, integrating visual information into planning or replanning during execution is necessary. This approach can significantly enhance the accuracy and feasibility of task planning, better addressing the challenges of real-world environments.
@@ VI-B2 Planning utilizing the visual information from embodied perception model
Based on the above discussion, it is particularly important to further integrate visual information into task planning (or replanning). In this process, object labels, locations, or descriptions provided by visual input can offer critical references for task decomposition and execution by LLMs. Through visual information, LLMs can more accurately identify target objects and obstacles in the current environment, thereby optimizing task steps or modifying subtask objectives. Some works use an object detector to query the objects present in the environment during task execution and feed this information back to the LLM, allowing it to modify unreasonable steps in the current plan [328, 326, 329]. RoboGPT considers the different names of similar objects within the same task, further improving the feasibility of replanning [10]. However, the information provided by labels is still too limited. Can further scene information be provided? SayPlan [330] proposes using hierarchical 3D scene graphs to represent the environment, effectively mitigating the challenges of task planning in large, multi-floor, and multi-room settings. Similarly, ConceptGraphs [331] also adopts 3D scene graphs to provide environmental information to LLMs. Compared to SayPlan, it offers more detailed open-world object detection and presents task planning in a code-based format, which is more efficient and better suited to the demands of complex tasks.
However, limited visual information can result in an agent’s inadequate understanding of its environment. While LLMs are provided with visual cues, they often fail to capture the environment’s complexity and dynamic changes, leading to misunderstandings and task failures. For example, if a towel is locked in a bathroom cabinet, the agent might repeatedly search the bathroom without considering this possibility [10]. To address this, more robust algorithms must be developed to integrate multiple sensory data, enhancing the agent’s environmental understanding. Additionally, leveraging historical data and contextual reasoning, even when visual information is limited, can aid the agent in making reasonable judgments and decisions. This approach of multimodal integration and context-based reasoning not only increases task execution success rates but also provides new perspectives for the advancement of embodied artificial intelligence.
@@ VI-B3 Planning utilizing the VLMs
Compared to converting environmental information into text using external visual models, VLM models can capture visual details in latent space, particularly contextual information that is difficult to represent with object labels. VLMs can discern rules underlying visual phenomena; for instance, even if a towel is not visible in the environment, it can be inferred that the towel might be stored in a cabinet. This process essentially demonstrates how abstract visual features and structured textual features can be more effectively aligned in latent space. In EmbodiedGPT [303], the Embodied-Former module aligns embodied, visual, and textual information, effectively considering the agent’s state and environmental information during task planning. Unlike EmbodiedGPT, which directly uses third-person perspective images, LEO [332] encodes 2D egocentric images and 3D scenes into visual tokens. This method effectively perceives 3D world information and executes tasks accordingly. Similarly, the EIF-Unknow model utilizes Semantic Feature Maps extracted from Voxel Features as visual tokens, which are input along with text tokens into a trained LLaVA model for task planning [333]. Furthermore, embodied multimodal foundation models, or VLA models, have been extensively trained with large datasets in studies like the RT series [11, 301], PaLM-E [300], and Matcha [334] to achieve alignment of visual and textual features in embodied scenarios.
@@ VI-C1 Action utilizing APIs
This hierarchical planning paradigm allows agents to focus on high-level task planning and decision-making while delegating specific action execution to policy models, simplifying the development process. The modularity of the task planner and action planner enables independent development, testing, and optimization, enhancing system flexibility and maintainability. This method allows the agent to adapt to various tasks and environments by invoking different action planners and facilitating modifications without requiring significant changes to the agent’s structure. However, invoking external policy models may introduce latency, potentially impacting response time and efficiency, especially in real-time tasks. The agent’s performance is critically dependent on the quality of the policy models. If the policy models are ineffective, the overall performance of the agent will suffer.
@@ VII-A Embodied World Model
LEGENT[351] 2024 3D Environments Generation
GRUtopia [352] 2024 Language Instruction, Task Generation, Navigation, Social Interaction, Manipulation
VII-A1 Generation-based Methods
@@ VII-A2 Prediction-based Methods
The prediction-based world model predicts and understands the environment by constructing and utilizing internal representations. By reconstructing corresponding features in the latent space based on provided conditions, it captures deeper semantics and associated world knowledge. This model maps input information to a latent space and operates within that space to extract and utilize high-level semantic information, thereby enabling the robots to perceive the essential representation of the world environment (e.g., I-JEPA[16], MC-JEPA[341], A-JEPA[342], Point-JEPA [353], IWM[343]) and more accurately perform embodied downstream tasks (e.g., iVideoGPT[344], IRASim[345]), STP[346], MuDreamer[347]). Compared to pixel-level information, latent feature can abstract and decouple various forms of knowledge, allowing the model to handle complex tasks and scenes more effectively and improve its generalization capability[354]. For instance, in spatiotemporal modeling, the world model needs to predict the post-interaction state of an object based on its current state and the nature of the interaction, combining this information with its internal knowledge. Specifically, an embodied world model generates dynamic predictions of the environment by integrating perceptual information and prior knowledge. This approach relies not only on sensory data but also on inherent world knowledge to infer and predict environmental changes, thereby producing more accurate spatiotemporal predictions[344, 346, 347]. This process considers both the current state of objects and their historical data and contextual information.
Figure 16: The illustration of demonstration data collection. The yellow box on the left features operational demonstrations of the Franka and WidowX robotic arms, while human demonstrations are shown in the blue box. In the middle, the yellow box showcases operational scenarios of the UR5e and Franka robotic arms in simulation environments, and the blue box displays labeled simulation data. The purple box on the right presents the data formats of these datasets.
@@ VII-A3 Knowledge-driven Methods
Knowledge-driven world models inject artificially constructed knowledge into the models, endowing them with world knowledge. This method has shown broad application potential in the field of embodied AI. For example, in the real2sim2real approach[356], real-world knowledge is used to build physics-compliant simulators, which are then used to train robots, enhancing model robustness and generalization capabilities. Additionally, artificially constructing common sense or physics-compliant knowledge and applying them to generative models or simulators is a common strategy (e.g., ElastoGen [349], One-2-3-45 [350], PLoT [348]). This approach imposes more physically accurate constraints on the model, enhancing its reliability and interpretability in generation tasks. These constraints ensure the model’s knowledge is both accurate and consistent, reducing uncertainty during training and application. Some approaches combine artificially created physical rules with LLMs or MLMs. By leveraging the commonsense capabilities of LLMs and MLMs, these approaches (e.g., Holodeck[70], LEGENT[351], GRUtopia [352]) generate diverse and semantically rich scenes through automatic spatial layout optimization. This greatly advances the development of general-purpose embodied agents by training them in novel and diverse environments.
@@ VII-C Embodied Control
Figure 18: Examples of embodied control for various locomotion modes, demonstrating the robots’ agile movement and interaction capabilities. The orange box on the left showcases the locomotion of quadruped robots (Unitree B1, ANYmal, Spot, Ghost Robotics Vision-60, MIT Cheetah) and one biped robot (ATRIAS). The red box on the middle displays motion scenes of the Atlas Robot climbing stairs, TELESAR V robot writing, Unitree H1 high-fiving, and the Boston robot leaping. The green box shows multi-robot collaborative control, finding path and planning path.
@@
Embodied control integrates RL and sim-to-real techniques to optimize strategies through environmental interaction, enabling the exploration of unknown domains, potentially surpassing human capabilities, and adapting to unstructured environments. While robots can mimic many human behaviors, effective task completion often requires RL training based on environmental feedback. The most challenging scenarios include contact-intensive tasks, where manipulation necessitates real-time adjustments based on feedback such as the state, deformation, material, and force of manipulated objects. In such cases, RL is indispensable. In the era of MLMs, these models possess a generalized understanding of scene semantics, providing robust reward functions for RL. Additionally, RL is crucial for aligning large models with their intended tasks. In the future, after pre-training and fine-tuning, RL is still required to align with the physical world, ensuring effective deployment in real-world environments.
@@ VIII Challenges and Future Directions
Efficient Utilization of Human Demonstration Data: Efficient utilization of human demonstration data involves leveraging the actions and behaviors demonstrated by humans to train and improve robotic systems. This process includes collecting, processing, and learning from large-scale, high-quality datasets where humans perform tasks that robots are intended to learn. Current work R3M [386] used action labels and human demonstration data to learn generalizable representations with high success rates, but the efficiency for complex tasks still needs improvement. Therefore, it is important to effectively utilize large amounts of unstructured, multi-label, and multi-modal human demonstration data combined with action label data to train embodied models that can learn various tasks in relatively short periods. By efficiently utilizing human demonstration data, robotic systems can achieve higher levels of performance and adaptability, making them more capable of performing complex tasks in dynamic environments.
@@
Long-Horizon Task Execution: Executing single instructions can often entail long-horizon tasks for robots, exemplified by commands like “clean the kitchen”, which involve activities such as rearranging objects, sweeping floors, wiping tables, and more. Accomplishing such tasks successfully necessitates the robot’s ability to plan and execute a sequence of low-level actions over extended time spans. While current high-level task planners have shown initial success, they often prove inadequate in diverse scenarios due to their lack of tuning for embodied tasks. Addressing this challenge requires the development of efficient planners equipped with robust perception capabilities and much commonsense knowledge.
Causal Relation Discovery: Existing data-driven embodied agents make decisions based on the intrinsic correlations within the data. However, this modeling approach does not allow the models to truly understand the causal relations between knowledge, behavior, and environment, resulting in biased strategies. This makes it difficult to ensure that they can operate in real-world environments in an interpretable, robust, and reliable manner. Therefore, it is important for embodied agents to be driven by world knowledge, capable of autonomous causal reasoning. By understanding the world through interaction and learning its workings via abductive reasoning, we can further enhance the adaptability, decision reliability, and generalization capabilities of multimodal embodied agents in complex real-world environments. For embodied tasks, it is necessary to establish spatial-temporal causal relations across modalities through interactive instructions and state predictions [387]. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [388], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied AI.